# MIVE: New Design and Benchmark for Multi-Instance Video Editing

### v1

@@ E.3 IPR Ablations
The final hyperparameter is the percentage of steps during which we apply our IPR (IPR steps percentage). Figure 27 illustrates the qualitative results of different configurations for the IPR steps percentage. Increasing the IPR steps percentage generally enhances editing faithfulness, e.g., better wood carvings on the table, improved rusty effects on the beaker, and more refined features on the robot’s head. However, excessively increasing the percentage may introduce artifacts, e.g., the appearance of noise on the glass table, due to the enhanced editing faithfulness, similar to the observations with $\lambda_{r}$. Our quantitative results in Tab. 9 align with these findings, showing that increasing the IPR steps percentage tends to improve editing faithfulness metrics (GTF, LTF, and IA). Interestingly, applying IPR to 100% of the sampling steps reduces editing faithfulness, likely due to the emergence of artifacts that disrupt details and negatively impact the editing faithfulness scores. Based on these observations, we choose to apply IPR to 80% of the sampling steps. Overall, the selection of these three hyperparameters in IPR is guided by the trade-off between editing faithfulness and the emergence of artifacts, which are the two key factors in cross-attention.
F Additional Details
F.1 Details on Long Video Editing
Our MIVE dataset comprises of several long videos, but directly editing them is constrained by GPU memory limitations. To address this, we leverage the keyframe selection algorithm proposed in FreSCo [62]. Using our MIVE framework, we edit the selected keyframes and employ the generative video interpolation method Ebsynth [24] to obtain the non-keyframes.
F.2 Running Time
Figure 28: Running time of our method. (a) Running time of our method vs number of video frames. (b) Running time of our method vs number of instances
Figure 28 illustrates the running time of our method on a $512\times 512$ video using a single 48GB-NVIDIA RTX A6000 GPU. The running time of our method is influenced by two main factors: (i) the number of frames and (ii) the number of instances to edit, both of which impact the running time linearly. On average, our method takes approximately 16.61 seconds per instance per frame. Regarding memory consumption, it is largely independent of the number of instances to edit, as this factor has only a minimal impact on memory usage. Instead, the memory consumption is primarily determined by the number of frames.
G Failure Case and Limitation
G.1 Failure Case
Figure 29: Failure case. Our method is less faithful when there is a large number of instances to edit. Note that other SOTA methods also fail to handle this large number of instances to edit.
We present a failure case in Fig. 29, where our MIVE struggles when faced with an excessive number of instances to edit. However, note that other SOTA methods, e.g., ControlVideo [70] and GAV [25], also struggle in this extreme scenario. One potential solution is to train an additional adapter on the LDM backbone that leverages the mask to joint modify features after the self-attention (similar to the approach in [55]) and after the cross-attention layers. This approach could better utilize the mask condition, thereby improving performance in scenarios involving a large number of instances in multi-instance video editing. However, this solution may reduce the flexibility of our method, as it would require re-training the adapter each time we switch to a different backbone.
G.2 Limitations
@@
Another limitation of our method is its running time. We introduce a new factor—the number of instances to edit—which impacts the running time linearly. Although the design of our method significantly improves editing faithfulness and reduces attention leakage, it also increases the running time. Consequently, the more instances that need to be edited, the longer the process takes. One potential solution to this issue is to replace our current sampling method with a faster approach, such as DPM-Solver [35] or DPM-Solver++ [34] and adjust our inversion process accordingly using DPM-Solver inversion [22].

### v2

@@ D.2 IPR Ablations
The first hyperparameter we determine is the percentage of steps during which we apply our IPR (IPR steps percentage). Figure 18 illustrates the qualitative results of different configurations for the IPR steps percentage. Increasing the IPR steps percentage generally enhances editing faithfulness, e.g., better facial texture details in the alien’s head. However, excessively increasing the percentage may lead to loss of structure. Our quantitative results in Tab. 9 align with these findings, showing that increasing the IPR steps percentage tends to improve editing faithfulness metrics (GTF, LTF, FA, and IA). Based on these observations, we choose to apply IPR to 10% of the sampling steps.
Figure 19 illustrates the results of varying the values of the hyperparameter $\lambda$. From these results, we observe that increasing $\lambda$ steers the editing more towards the edit caption. For example, the alien’s face is slowly losing the man’s features as we increase $\lambda$. However, excessively increasing $\lambda$ may cause oversaturation as observed in the alien’s face and clothes. In terms of quantitative performance, both $\lambda=0.5$ and $\lambda=0.6$ yield comparable results. However, due to the potential risk of oversaturation, we use $\lambda=0.5$ as our default.
For the next hyperparameter, $\lambda_{r}$, in our IPR, the qualitative results are shown in Fig. 20. We observe that increasing $\lambda_{r}$ tends to increase details but introduces artifacts. For example,while the alien’s face becomes more detailed, it also exhibits noticeable noise, partially reverting to its original human form. These observations are further supported by the quantitative comparison in Tab. 9, where increasing $\lambda_{r}$ reduces faithfulness scores (GTF, FA, LTF, and IA). Based on these findings, we select $\lambda_{r}=0.5$ to achieve a balance between overall editing faithfulness and the minimization of artifacts.
Overall, the selection of these three hyperparameters in IPR is guided by the trade-off between editing faithfulness, emergence of artifacts, and oversaturation.
D.3 DMS Ablations
Figure 21: Ablation study on DMS: (a) Ablation on the number of SNS steps. Increasing the number of SNS steps reduces artifacts in the background and improves local faithfulness and background preservation scores. Setting SNS steps to the first 40% of the sampling process achieves a balanced trade-off between qualitative and quantitative performance. See Tab. 10-(a) for quantitative results.
Figure 22: Ablation study on DMS: (b) Ablation on the number of re-inversion steps $l$. Increasing the number of re-inversion steps enhances the global scores as well as the local temporal consistency performance of PRIMEdit. Quantitative results are provided in Tab. 10-(b). However, we set the number of re-inversion steps to $l=2$ to avoid blurring artifacts shown in (4).
Method Global Scores Local Scores Leakage Scores
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
(a) Ablation on the number of SNS steps
(1) No SNS (Pure PNS) 0.9670 0.2638 0.8008 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
(2) First 10% 0.9667 0.2634 0.7855 0.9541 0.2047 0.5356 0.6704 0.9009 0.0565
(3) First 20% 0.9663 0.2635 0.7780 0.9535 0.2050 0.5409 0.6739 0.9021 0.0551
(4) First 30% 0.9662 0.2631 0.7687 0.9534 0.2052 0.5390 0.6718 0.9031 0.0537
(5) First 40% (Ours) 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
(b) Ablation on the number of re-inversion steps
(1) l = 0 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
(2) l = 1 0.9671 0.2629 0.7710 0.9544 0.2050 0.5319 0.6713 0.9029 0.0549
(3) l = 2 (Ours) 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
(4) l = 3 0.9682 0.2635 0.7734 0.9560 0.2045 0.5342 0.6714 0.8976 0.0637
We conduct ablation studies on the hyperparameters of our Disentangled Multi-instance Sampling (DMS) technique. The quantitative results are summarized in Tab. 10, while the qualitative comparisons are shown in Figures 21 and 22.
First, we determine the optimal number of SNS steps during sampling. Without SNS, the framework struggles with background preservation and introduces artifacts (Fig. 21-(1)). As the number of SNS steps increases, artifacts gradually decrease (Fig. 21-(2) to (5)), leading to improved local faithfulness and background preservation scores, though with a slight decrease in temporal consistency. Based on our findings, setting SNS steps to the first 40% of the sampling process achieves a balanced trade-off between qualitative and quantitative performance (Tab. 10-(a)).
Finally, we conduct experiments by introducing re-inversion between the SNS and PNS steps. Increasing the number of re-inversion steps $l$ after SNS improves global metrics as well as the local temporal consistency (Tab. 10-(b)). However, excessively increasing $l$ leads to blurring and artifacts Fig. 22-(4). Ultimately, we set $l=2$ to balance the overall qualitative appearance and quantitative performance.
E Qualitative Results on DAVIS Dataset
Figure 23: Qualitative results on four DAVIS Dataset videos. Our PRIMEdit maintains faithful and temporally consistent edits, particularly in cases with large object motions (rows 3 and 4).
Fig. 23 presents qualitative results on four videos from the DAVIS dataset, demonstrating the effectiveness of our PRIMEdit in maintaining faithful and temporally consistent edits. Our approach preserves object details, even in challenging scenarios involving large object motions (rows 3 and 4). These results highlight the robustness of our method in achieving high-quality video edits with strong temporal coherence.
F Efficiency Comparison with VideoGrain
Method Time (sec) $\downarrow$ Memory (GB) $\downarrow$
Single Instance, 15 frames
VideoGrain [65] 328.28 21.96
PRIMEdit (Ours) 143.27 15.28
Single Instance, 32 frames
VideoGrain [65] 1181.41 41.90
PRIMEdit (Ours) 295.43 24.24
3 Instances, 15 frames
VideoGrain [65] 331.39 22.86
PRIMEdit (Ours) 225.71 15.33
3 Instances, 32 frames
VideoGrain [65] 1181.08 46.06
PRIMEdit (Ours) 465.06 24.3
Table 11: Runtime and GPU memory usage comparison between our PRIMEdit and VideoGrain [65]. The best and second best scores are shown in red and blue, respectively.
We compare runtime and GPU memory usage against VideoGrain [65] in Tab. 11, using a single RTX A6000 GPU for editing under four scenarios: (i) a single instance over 15 frames, (ii) a single instance over 32 frames, (iii) three instances over 15 frames, and (iv) three instances over 32 frames. As shown in Tab. 11, our PRIMEdit framework is both faster and more memory-efficient for multi-instance video editing.
G Limitations